# KCL-PH-TH/2025-33

### v2

@@ 1 Introduction
The standard cosmological narrative begins with a period of slow-roll inflation [30, 35, 2], during which the dynamics are governed by a scalar field—the inflaton—rolling down a nearly flat potential. This phase provides solutions to the horizon and flatness problems, while also stretching quantum fluctuations of the inflaton field, which later seed the formation of large-scale structures. After inflation ends, it is usually assumed that the inflationary sector reheats the universe by transferring most of its energy to Standard Model particles. This marks the beginning of the Hot Big Bang phase, initially characterized by a radiation-dominated plasma with an equation-of-state parameter $w=1/3$, which eventually transitions to matter domination with $w=0$, as the universe cools. Interestingly, there are no strong constraints on the equation of state of the universe during the period between the end of inflation and the onset of reheating. This opens up the possibility for a variety of cosmological scenarios.
One such possibility is that the post-inflationary–but pre-reheating–epoch is a phase during which the universe’s dynamics are dominated by the kinetic energy of a rolling scalar field, which may be the inflaton or another scalar. This phase, referred to as kination, was first proposed in [34], characterised by equation-of-state parameter $w=1$. For a more comprehensive review of kination we refer the reader to [29]. Kination is especially relevant in the context of string theory, where ultra-light moduli fields can dominate early-universe dynamics [15]. String cosmology models involving either the inflaton or other moduli may naturally give rise to kination if the relevant modulus potential features a steep region at the end of inflation–where the potential decreases exponentially, enabling the field to acquire significant kinetic energy–followed by a flat region where the field eventually slows down. What happens afterward depends on the exact shape of the potential.
Suppose the potential is of a runaway form, the field slows down via Hubble friction, possibly initiating a secondary period of slow-roll inflation. It is worth noting that this feature is common in quintessential inflation models [38, 37, 19, 11, 22, 23] and more generally, in non-oscillatory inflation scenarios [25, 33], where the inflaton field undergoes $(i)$ a slow-roll phase in the early universe, $(ii)$ a phase of acceleration as the potential decreases sharply and the universe enters into a kination era, and $(iii)$ a second phase of slow-roll during which the field decelerates to eventually either behave like a cosmological constant or track the universe’s background energy density at late times [43, 41, 25, 33, 5]. On the other hand, if the potential features a local minimum, the field may eventually settle into it, acquiring a finite mass and oscillating before decaying to reheat the universe, as occurs in the so-called Large Volume Scenarios (LVS) [10, 18, 5].
@@
In [5, eröncel2025universalbounddurationkination, 36], the fate of small scalar perturbations is explored in the context of an initially kinetic-energy-dominated background where the gravitational backreaction is ignored. Perturbation theory has provided valuable insights into this regime; however, its main limitation is that it cannot capture strong gravitational effects or backreaction.
@@
In this work, we are instead interested in the evolution of kination in the presence of non-linear perturbations where strong gravitational effects and backreaction can play an important role. This becomes particularly important during kination, where even initially small scalar field perturbations—especially those of large super-horizon wavelength—can be significantly amplified. As these modes re-enter the Hubble horizon, they may dominate the local dynamics, inducing sharp local gradient inhomogeneities that strongly backreact on the background geometry. Consequently, the study of perturbations during kination transitions into a strong-gravity problem. In scenarios where the backreaction becomes sufficiently strong, gravitational collapse may occur, potentially leading to black hole formation. In this context, while perturbation theory provides essential guidance, this regime necessitates the use of tools from numerical relativity.
@@
We find that for sub-horizon perturbations, even for relatively large initial amplitudes, gravitational backreaction remains weak and does not induce collapse. In this case, scalar perturbations on average evolve like radiation for scales much larger than the perturbation scales, in agreement with perturbative predictions. Gravitational backreaction only becomes significant with initial perturbation energy density much larger than the kination background energy density. In this scenario, the notion of a kination background is no longer clear. On the other hand, for linear super-horizon perturbations, gravitational collapse becomes significantly more likely after horizon re-entry. We provide a quantitative analysis of the critical initial overdensity required for collapse and compare it with perturbative expectations. Our results indicate that black hole formation is generally easier during kination than in perturbative analysis. This has important implications for the shape and amplitude of the power spectrum necessary to produce a significant primordial black hole abundance in the early universe.
@@
This paper is organised as follows: In Section 2 we provide an introductory framework on kination and our numerical simulation setups, in Section 3 we present our numerical results for sub-horizon and super-horizon perturbations respectively, focusing on the conditions for black hole formation in the super-horizon case. In Section 4 we discuss cosmological implications for the production of primordial black holes in cosmology, and we conclude in Section 5. In this work, we use the non-reduced Planck mass, $m_{P}^{2}\equiv\hbar c/G=G^{-1}$ with $\hbar=c=1$ for theoretical analysis, and we use geometric units with $G=c=1$ for simulation works.
@@ 3.2 Super-Horizon Gradient Perturbations
For inhomogeneities with characteristic wavelength $\lambda_{0}>H_{0}^{-1}$, linear perturbation theory indicates that the perturbation energy density of super-horizon modes is approximately given by the gradient term and scales as $\rho_{\delta\phi}\sim\rho_{\nabla}\sim a^{-2}$. Thus, the density contrast grows as $a^{4}$ compared to the kination background and the modes behave like scalar curvature perturbations. However, in practice, as shown in the previous section, the energy density of the perturbations consists of both the gradient term and the induced kinetic term, but the superhorizon kinetic part according to the perturbative analysis tracks the background and decays as $\rho_{\delta K}\sim a^{-6}$, therefore, very quickly dilutes, we
We begin by confirming this in the linear limit and simulate a scalar field with a homogeneous kinetic energy profile and an inhomogeneity characterised by an initial wavelength of $\lambda_{0}=2\times 10^{4}H_{0}^{-1}$. The amplitude of this inhomogeneity is chosen by setting the initial density contrast to $\delta_{0}=4\times 10^{-10}$. The results of the simulation are shown in Figure 2. We find that the volume-averaged perturbation energy density $\langle\rho_{\delta\phi}\rangle$ very quickly tracks the gradient energy density once the system relaxes from the initial conditions. This happens within approximately one e-fold. We also find that the variance of the local expansion parameter $\langle\delta K\rangle/\langle K\rangle\ll 1$, indicating that the simulation domain remains highly homogeneous and the spacetime closely follows the FLRW metric. Thus, the perturbation energy density then evolves as $\langle\rho_{\delta\phi}\rangle\sim\langle\rho_{\nabla}\rangle\sim a^{-2}$.
@@
This estimate assumes a consistent and well-defined notion of the scale factor throughout the simulation domain and evolution. To ensure this, we verify that the domain remains well-described by an FLRW spacetime. Following the methods of [27, 7] we evaluate the variance, $\sigma^{2}$, of the local expansion parameter. In scenarios where black hole formation is expected, we additionally monitor the variance of the conformal factor and compute their kurtoses, $g_{2}$ [^2] We use the natural estimator definition of $g_{2}=\frac{\frac{1}{n}\sum_{i}^{n}(x_{i}-\bar{x})^{4}}{\sigma^{4}}-3,$ where $\sigma$ is the standard deviation. Under this definition, $g_{2}=-3$ indicates absolute uniformity, $g_{2}=0$ indicates a normal distribution, and $g_{2}>0$ indicates a sharp peak in distribution., to assess deviation from FLRW behaviour near the onset of gravitational collapse. In the simulation shown in Figure 3, we observe that both $\sigma^{2}_{K}$ and $\sigma^{2}_{\chi}$ remain in the range $\mathcal{O}(10^{-3}\sim 10^{-5})$ throughout the evolution. We also find that $g_{2,\chi}\sim g_{2,K}\sim-1$ prior to the black hole formation showing platykurtic(or near uniformity) distribution. These results confirm that the spacetime remains a good approximation to FLRW prior to the black hole formation.
In the simulation example shown in Figure 3, we estimate the horizon re-entry to occur at $\log(a)_{\mathrm{re-entry}}\approx 1.5$ with $\langle\rho_{\delta\phi}\rangle\lesssim\overline{\rho}$ and black hole formation at $\log(a)_{\mathrm{BH}}\approx 1.9$ as indicated by the discontinuity in the energy density profiles. The first detection of an apparent horizon and the estimated mode horizon re-entry are separated by approximately 0.5 e-folds. We observed that this behaviour is consistent across all super-horizon modes collapses we have tested.
Given the relative growth of the perturbation, the simulation domain deviates from pure FLRW spacetime, making the precise moment of horizon re-entry difficult to define. We define $\delta_{c,\text{ effective}}$ as the effective critical overdensity, extrapolated from the simulation results under the assumptions that the kinetic energy density always dominates the universe’s evolution prior to horizon re-entry, and that the perturbation grows relative to the kination background remains as $a^{4}$
@@
Note that this quantity is defined for the sake of comparing our results to the perturbative case, but does not correspond to a quantity that is actually measurable numerically, unlike $\delta_{0,\text{crit}}$, which is the physical and measurable quantity in our approach. As mentioned above, we aim to investigate whether the critical overdensity is robust for super-horizon modes that exhibit non-linear behaviour before horizon re-entry. We simulated a range of super-horizon modes with initial wavelengths spanning from $5H_{0}^{-1}$ to $2000H_{0}^{-1}$ and an initial density contrast $\delta_{0}$ ranging from $\mathcal{O}(10^{-2}\sim 10^{-9})$. We present the results in Figure 4.
@@
Note that the presented values do not represent the critical overdensity for primordial black hole formation at horizon re-entry. Instead, they serve as a comparison to the theoretical predictions if one expects the super-horizon modes to stay frozen as in the perturbative limit, until re-entry, i.e $\delta\sim a^{4}$. Under this assumption, these are the values that correspond to the critical initial conditions that lead to primordial black holes. We detail the cosmological implications in the following section.
@@
    As discussed above, we observe the point of horizon re-entry occurs at $a_{\text{re-entry}}\approx\sqrt{\lambda_{0}H_{0}}$ with black hole formation typically taking place about $\sim 0.5$ e-folds later. In this work, the largest scale factor we were able to probe is $a\sim 80$. To investigate longer super-horizon modes that would re-enter the horizon during a prolonged kination era lasting up to $N_{\text{kin}}\sim 10$, the simulation would need to reach scale factors of $\mathcal{O}(10^{3}\sim 10^{4})$. However, achieving such large scale factors is currently infeasible with the numerical setup employed in this study. Numerical stability constrains the maximum allowed time step, and reaching significantly larger values of $a$ leads to an accumulation of numerical errors and requires computational resources that grow exponentially with simulation duration.
@@
    The gauge choice we have imposed in Eq. (2.17) ensures that the presence of a small perturbation does not induce strong gravitational backreaction until close to horizon re-entry. Thus the simulation domain is effectively a uniformly expanding spacetime with a decaying expansion rate. In such domains, we observe the local expansion parameter is $K\lesssim 0$, with regions fluctuating slightly above and below the average. While Eq. (2.17) is a more cosmology-friendly gauge [7], the prolonged oscillation periods for longer super-horizon modes gradually drive the lapse value across the entire hyperslice downward. As detailed in 2.3, the lapse function is initially set to 1 everywhere on the hyperslice. The combined effect of the dynamics and this gauge choice causes the lapse function to drop to $\mathcal{O}(10^{-1})$ across the domain, thereby increasing both the simulation time and computational cost.
@@ 3.3 Super-Horizon Kinetic Perturbations
For completeness, we also investigated the effects of a pure kinetic super-horizon perturbation in this regime. For such a mode, perturbation theory suggests $\rho_{\delta K}\sim a^{-6}$ (see Appendix. A). However, as we have seen in the previous section that the super-horizon gradient perturbation redshifts as $a^{-2}$, the induced super-horizon gradient mode would eventually become the dominant contribution of the scalar perturbation. We present the result from the simulation with the same $\lambda_{0}$ and $\delta_{0}$ as in Figure 3. We observe that initially, the kinetic perturbation drops drastically to induce gradient energy until they are of the same order. The total perturbation energy density then behaves roughly as $a^{-2}$ before the estimated re-entry point. However, since the gradient energy density is induced as a later time than in the simulation of Figure 3 and at a lower order, it does not have enough time to grow to the same non-linear regime as in the previous simulation, and thus do not undergo gravitational collapse. After the estimated re-entry at around $\log(a)\sim 1.5$, it redshifts as $a^{-4}$, which again confirms the result for the sub-horizon case.
Figure 6: Simulation results for super-horizon kinetic perturbation with initial conditions of $\lambda_{0}=20H_{0}^{-1}$, $\delta_{0}=5\times 10^{-4}$. Upper panel: the evolution of volume-averaged energy densities of the background, gradient, kinetic, and perturbations, Lower panel: The background, gradient, kinetic, and perturbation energy densities as a power law of the $\log$ of the volume-averaged scale factor $a=\langle\chi\rangle^{-1/2}$.

### v3

@@ 1 Introduction
The standard cosmological narrative begins with a period of slow-roll inflation [30, 35, 2], during which the dynamics are governed by a scalar field—the inflaton—rolling down a nearly flat potential. This phase provides solutions to the horizon and flatness problems while stretching quantum fluctuations of the inflaton field that later seed the formation of large-scale structures. After inflation ends, it is typically assumed that the inflationary sector reheats the universe by transferring most of its energy to Standard Model particles. This marks the beginning of the Hot Big Bang phase, initially characterized by a radiation-dominated plasma with equation-of-state parameter $w=1/3$, which eventually transitions to matter domination with $w=0$ as the universe cools.
Interestingly, there are no strong observational constraints on the equation of state of the universe during the period between the end of inflation and the onset of reheating. This opens up the possibility for a variety of cosmological scenarios. One such possibility is that the post-inflationary–but pre-reheating–epoch is dominated by the kinetic energy of a rolling scalar field, which may either be the inflaton or another scalar degree of freedom. This phase, known as kination, was first proposed in [34], characterised by equation-of-state parameter $w=1$. For a more comprehensive review of kination, see [29].
Kination is especially relevant in the context of string theory, where ultra-light moduli fields can dominate early-universe dynamics [15]. String cosmology models involving either the inflaton or other moduli may naturally give rise to kination if the relevant modulus potential features a steep region at the end of inflation–where the potential decreases exponentially, enabling the field to acquire significant kinetic energy–followed by a flat region where the field eventually slows down. What happens afterward depends on the exact shape of the potential.
If the potential exhibits runaway behaviour, the field slows down due to the Hubble friction and may even initiate a secondary period of slow-roll inflation. This feature is common in quintessential inflation models [38, 37, 19, 11, 22, 23] and more generally, in non-oscillatory inflation scenarios [25, 33], where the inflaton field undergoes $(i)$ a slow-roll phase in the early universe, $(ii)$ a phase of acceleration as the potential decreases sharply and the universe enters a kination era, and $(iii)$ a second phase of slow-roll during which the field decelerates to eventually either behave as an effective cosmological constant or track the universe’s background energy density at late times [43, 41, 25, 33, 5]. Alternatively, if the potential features a local minimum, the field may eventually settle into it, acquiring a finite mass and oscillating before decaying to reheat the universe, as occurs in Large Volume Scenarios (LVS) [10, 18, 5].
@@
In [5, eröncel2025universalbounddurationkination, 36], the fate of small scalar perturbations is explored in the context of an initially kinetic-energy-dominated background where the gravitational backreaction is neglected. Perturbation theory has provided valuable insights into this regime; however, its main limitation is that it cannot capture strong gravitational effects or non-linear backreaction.
@@
In this work, we instead investigate the evolution of kination in the presence of non-linear perturbations where strong gravitational effects and backreaction can play an important role. This becomes particularly important during kination, where even initially small scalar field perturbations—especially those of large super-horizon wavelength—can undergo signification amplification. As these modes re-enter the Hubble horizon, they may dominate the local dynamics, inducing sharp local gradient inhomogeneities that strongly backreact on the geometry. Consequently, the study of perturbations during kination transitions into a genuine strong-gravity problem. In scenarios where the backreaction becomes sufficiently strong, gravitational collapse may occur, potentially leading to black hole formation. In this context, while perturbation theory provides essential guidance, this regime requires numerical relativity.
@@
We find that for sub-horizon perturbations, even for relatively large initial amplitudes, gravitational backreaction remains weak and does not induce collapse. In this case, scalar perturbations on average evolve like radiation for scales much larger than the perturbation scales, in agreement with perturbative predictions. Gravitational backreaction only becomes significant with initial perturbation energy density much larger than the kination background energy density. In this scenario, the notion of a kination background ceases to be well defined. In contrast, for initially linear super-horizon perturbations, gravitational collapse becomes substantially more likely after horizon re-entry. We provide a quantitative analysis of the critical initial overdensity required for collapse and compare it with perturbative expectations. Our results indicate that black hole formation is generally easier during kination than in perturbative analysis. This has important implications for the shape and amplitude of the power spectrum necessary to produce a significant primordial black hole abundance in the early universe.
@@
This paper is organised as follows: In Section 2 we provide an introductory framework on kination and our numerical simulation setups, in Section 3 we present our numerical results for sub-horizon and super-horizon perturbations respectively, focusing on the conditions for black hole formation in the super-horizon case. In Section 4 we discuss cosmological implications for the production of primordial black holes in cosmology, and we conclude in Section 5.
In this work, we use the non-reduced Planck mass, $m_{P}^{2}\equiv\hbar c/G=G^{-1}$ with $\hbar=c=1$ for theoretical analysis, and we use geometric units with $G=c=1$ for simulation works.
@@ 3.2 Super-Horizon Gradient Perturbations
For inhomogeneities with characteristic wavelength $\lambda_{0}>H_{0}^{-1}$, linear perturbation theory indicates that the perturbation energy density of super-horizon modes is approximately given by the gradient term and scales as $\rho_{\delta\phi}\sim\rho_{\nabla}\sim a^{-2}$. Thus, the density contrast grows as $a^{4}$ compared to the kination background and the modes behave like scalar curvature perturbations. We begin by confirming this in the linear limit and simulate a scalar field with a homogeneous kinetic energy profile and an inhomogeneity characterised by an initial wavelength of $\lambda_{0}=1\times 10^{3}H_{0}^{-1}$. The amplitude of this inhomogeneity is chosen by setting the initial density contrast to $\delta_{0}=1\times 10^{-10}$. The results of the simulation are shown in Figure 2. We find that the volume-averaged perturbation energy density $\langle\rho_{\delta\phi}\rangle$ very quickly tracks the gradient energy density once the system relaxes from the initial conditions. This happens within approximately one e-fold. We also find that the variance of the local expansion parameter $\langle\delta K\rangle/\langle K\rangle\ll 1$, indicating that the simulation domain remains highly homogeneous and the spacetime closely follows the FLRW metric. Thus, the perturbation energy density then evolves as $\langle\rho_{\delta\phi}\rangle\sim\langle\rho_{\nabla}\rangle\sim a^{-2}$.
@@
This estimate assumes a consistent and well-defined notion of the scale factor throughout the simulation domain and evolution. To ensure this, we verify that most of the simulation domain remains well-described by an FLRW spacetime. Following the methods of [27, 7] we evaluate the variance, $\sigma^{2}$, of the local expansion parameter. In scenarios where black hole formation is expected, we additionally monitor the variance of the conformal factor and compute their kurtoses, $g_{2}$ [^2] We use the natural estimator definition of $g_{2}=\frac{\frac{1}{n}\sum_{i}^{n}(x_{i}-\bar{x})^{4}}{\sigma^{4}}-3,$ where $\sigma$ is the standard deviation. Under this definition, $g_{2}=-3$ indicates absolute uniformity, $g_{2}=0$ indicates a normal distribution, and $g_{2}>0$ indicates a sharp peak in distribution., to assess deviation from FLRW behaviour near the onset of gravitational collapse. In the simulation shown in Figure 3, we observe that both $\sigma^{2}_{K}$ and $\sigma^{2}_{\chi}$ remain in the range $\mathcal{O}(10^{-3}\sim 10^{-5})$ throughout the evolution. We also find that $g_{2,\chi}\sim g_{2,K}\sim-1$ prior to the black hole formation showing platykurtic(or near uniformity) distribution. These results confirm that the spacetime remains a good approximation to FLRW prior to the black hole formation.
In the simulation example shown in Figure 3, we estimate the horizon re-entry to occur at $\log(a)_{\mathrm{re-entry}}\approx 1.5$ with $\langle\rho_{\delta\phi}\rangle\lesssim\overline{\rho}$ and black hole formation at $\log(a)_{\mathrm{BH}}\approx 1.9$ as indicated by the discontinuity in the energy density profiles. The first detection of an apparent horizon and the estimated mode horizon re-entry are separated by approximately 0.5 e-folds. We observed that this behaviour is consistent across all super-horizon modes collapse we have tested. We note that while we can still attribute FLRW cosmological terms to the simulation, the simulation domain does deviate from a pure FLRW spacetime, making the precise moment of horizon re-entry difficult to define.
We define $\delta_{c,\text{ effective}}$ as the effective critical overdensity, extrapolated from the simulation results under the assumptions that the kinetic energy density always dominates the universe’s evolution prior to horizon re-entry, and that the perturbation grows relative to the kination background remains as $a^{4}$
@@
Note that this quantity is defined for the sake of comparing our results to the perturbative case, but does not correspond to a quantity that is actually measurable numerically, unlike $\delta_{0,\text{crit}}$, which is the physical and measurable quantity in our approach. As mentioned above, we aim to investigate whether the critical overdensity is robust for super-horizon modes that exhibit non-linear behaviour before horizon re-entry. We simulated a range of super-horizon modes with initial wavelengths spanning from $5H_{0}^{-1}$ to $2000H_{0}^{-1}$ and an initial density contrast $\delta_{0}$ ranging from $\mathcal{O}(10^{-2}\sim 10^{-10})$. We present the results in Figure 4.
@@
We emphasise again that the presented values do not represent the critical overdensity for primordial black hole formation at horizon re-entry. Instead, they serve as a comparison to the theoretical predictions if one expects the super-horizon modes to stay frozen as in the perturbative limit, until re-entry, i.e $\delta\sim a^{4}$. Under this assumption, these are the values that correspond to the critical initial conditions that lead to primordial black holes. We detail the cosmological implications in the next section.
@@
    As discussed above, we estimate the point of horizon re-entry occurs at $a_{\text{re-entry}}\approx\sqrt{\lambda_{0}H_{0}}$ with black hole formation typically taking place about $\sim 0.5$ e-folds later. In this work, the largest scale factor we were able to probe is $a\sim 80$. To investigate larger super-horizon modes that would re-enter the horizon during a prolonged kination era lasting up to $N_{\text{kin}}\sim 10$, the simulation would need to reach scale factors of $\mathcal{O}(10^{3}\sim 10^{4})$. However, achieving such large scale factors is currently infeasible with the numerical setup employed in this study. Numerical stability constrains the maximum allowed time step, and reaching significantly larger values of $a$ leads to an accumulation of numerical errors and requires computational resources that grow exponentially with simulation duration.
@@
    The gauge choice we have imposed in Eq. (2.17) ensures that the presence of a small perturbation does not induce strong gravitational backreaction until close to horizon re-entry. Thus the simulation domain is effectively a uniformly expanding spacetime with a decaying expansion rate. In such domains, we observe the local expansion parameter is $K\lesssim 0$, with regions fluctuating slightly above and below the average. While Eq. (2.17) is a more cosmology-friendly gauge [7], the prolonged oscillation periods for longer super-horizon modes gradually drive the lapse value across the entire hyperslice downward. As detailed in 2.3, the lapse function is initially set to 1 everywhere on the hyperslice. The combined effect of the dynamics and this gauge choice causes the lapse function to slowly drop to $\mathcal{O}(10^{-1})$ across the domain, thereby increasing both the simulation time and computational cost.
@@ 3.3 Super-Horizon Kinetic Perturbations
For completeness, we also investigated the effects of a pure kinetic super-horizon perturbation in this regime. For such a mode, perturbation theory suggests $\rho_{\delta K}\sim a^{-6}$ (see Appendix. A). However, as we have seen in the previous section that the super-horizon gradient perturbation redshifts as $a^{-2}$, the induced super-horizon gradient mode would shortly become the dominant contribution of the scalar perturbation. We present the result from the simulation with the same $\lambda_{0}$ and $\delta_{0}$ as in Figure 3. We observe that initially, the kinetic perturbation drops drastically to induce gradient energy until they are of the same order. The total perturbation energy density then behaves roughly as $a^{-2}$ before the estimated re-entry point. However, since the gradient energy density is induced at a later time than in the simulation of Figure 3 and at a lower order, it does not have enough time to grow to the same non-linear regime as in the previous simulation, and thus do not undergo gravitational collapse. After the estimated re-entry at around $\log(a)\sim 1.5$, the total scalar perturbation energy density redshifts as $a^{-4}$, which again confirms the result for the sub-horizon case.
Figure 6: Simulation results for super-horizon kinetic perturbation with initial conditions of $\lambda_{0}=20H_{0}^{-1}$, $\delta_{0}=5\times 10^{-4}$. Upper panel: the evolution of volume-averaged energy densities of the background, gradient, kinetic, and perturbations, Lower panel: The background, gradient, kinetic, and perturbation energy densities as a power law of the $\log$ of the volume-averaged scale factor $a\approx\langle\chi\rangle^{-1/2}$.